### v1

# Non-monotonic constitutive curves and shear banding in dry and wet granular flows

Abstract
We use particle simulations to map comprehensively the shear rheology of dry and wet granular matter, in both fixed pressure and fixed volume protocols. At fixed pressure we find non-monotonic constitutive curves that are shear thinning, whereas at fixed volume we find non-monotonic constitutive curves that are shear thickening. We show that the presence of one non-monotonicity does not imply the other. Instead, there exists a signature in the volume fraction measured under fixed pressure that, when present, ensures non-monotonic constitutive curves at fixed volume. In the context of dry granular flow we show that gradient and vorticity bands arise under fixed pressure and volume respectively, as implied by the constitutive curves. For wet systems our results are consistent with a recent experimental observation of shear thinning at fixed pressure. We furthermore predict discontinuous shear thickening in the absence of critical load friction.
Dense granular packings, both dry and suspended in liquid, are among the most abundant materials on earth. They are relevant to manifold geophysical phenomena, e.g., landslides and debris flows [1, 2], and to industrial processes such as paste extrusion [3, 4]. Understanding their deformation and flow properties is thus of major practical importance, and is of fundamental interest in statistical physics, fluid mechanics and rheology [5, 6, 7, 8].
For any complex fluid, a key rheological fingerprint is the constitutive relation of shear stress $\sigma_{xy}$ as a function of shear rate $\dot{\gamma}$ in stationary homogeneous shear. For granular materials, there exist two paradigmatic protocols for characterising this relation. In the first (“fixed volume”) [9, 10, 11, 12, 13, 14], one fixes the sample volume $V$ and measures $\sigma_{xy}$ (and sometimes the normal stress or normal stress differences) as a function of $\dot{\gamma}$ at a given particle volume fraction $\phi$. In the second (“fixed pressure”), one fixes the external pressure (usually in fact the normal particle stress ${\sigma_{yy}}$; but see [15]) and allows $\phi$ to vary while measuring $\sigma_{xy}$ as a function of $\dot{\gamma}$.
In the fixed-$V$ protocol, dry systems display distinct quasistatic, intermediate and inertial flow regimes dependent on $\dot{\gamma}$ and $\phi$ [11]. In suspensions the inertial regime is replaced by a viscous regime [16, 17]. Hysteresis is often observed close to the jamming volume fraction $\phi_{\rm m}$ that marks the transition between quasistatic and inertial or viscous regimes, suggesting that the constitutive curve $\sigma_{xy}(\dot{\gamma})$ has a shear-thickening, S-shaped non-monotonicity. In the unstable region ${\sigma_{xy}}$ can adopt multiple values at a single ${\dot{\gamma}}$ [18, 19, 20], implying a predisposition to vorticity banding [21].
In the fixed-$\sigma_{yy}$ protocol, experimental data for the macroscopic friction $\mu={\sigma_{xy}}/{\sigma_{yy}}$ across a range of scaled shear rates (“inertial number”) $I={\dot{\gamma}}a/\sqrt{{\sigma_{yy}}/\rho}$ [22] suggests universal constitutive relations $\mu(I)$ and $\phi(I)$ in dry systems [23, 24]. Here $a$ and $\rho$ are the particle radius and density. Recent innovations enabling fixed-${\sigma_{yy}}$ measurements in suspensions [25, 26] remarkably suggest analogous relations $\mu(J)$ and $\phi(J)$, with $J=\eta{\dot{\gamma}}/{\sigma_{yy}}$ now the “viscous number”, where $\eta$ is the solvent viscosity [25, 27]. In slow shear (small $I,J$), $\mu$ was shown to decrease with increasing $I$ in dry simulations [28] and experiments [29] (see also [30, 31, 32, 33]), and likewise suggested to decrease with increasing $J$ in experiments on wet systems [34], before increasing at large $I,J$. This gives a shear-thinning non-monotonicity, in which $\dot{\gamma}$ can adopt multiple values at a single $\mu$, implying a predisposition to gradient banding.
Figure 1: Non-monotonic constitutive curves of dry granular material at fixed $\Sigma_{yy}$ (a)-(e) and fixed $V$ (f)-(h). Shown are (a) $\mu(I)$ and (c) $\phi(I)$, respectively zoomed in (b), (d). Each curve of color blue to red has increasing $\dot{\gamma}_{d}$ left to right at a fixed $\Sigma_{yy}$ shown by the legend in (a). Purple line in (d) connects points with ${\dot{\gamma}}_{d}=4\times 10^{-6}$ and varying ${\Sigma_{yy}}$. Broad grey lines show fits of [22] (a) and [25] (c). Our fixed-$\Sigma_{yy}$ feedback algorithm maintains a set-point $\Sigma_{yy}^{\mathrm{SP}}=1.6\times 10^{-3}$ over time (e). In (f) are fixed-$V$ time series of $\Sigma_{xy}$ at $\dot{\gamma}_{d}=10^{-5}$ for various $\phi$. In (g),(h) we first replot fixed-$\Sigma_{yy}$ data from (a)-(d) as faded lines blue to red, then connect points of equal $\phi$ on these curves by lines grey to black. This gives fixed-$V$ constitutive curves, reconstructed from the fixed-$\Sigma_{yy}$ data. Shown by darker red, green and blue lines in (g),(h) are constitutive curves measured actually at fixed $V$, by binning time series of $\Sigma_{xy}$ (shown in (f)) and $\Sigma_{yy}$ as described in the text. Shaded boxes in (b),(d),(g),(h) cover regimes that have $\frac{\partial\mu}{\partial I}|_{\Sigma_{yy}}<0$ (pink) and $\frac{\partial\phi}{\partial I}|_{\dot{\gamma}_{d}}>0$ (green), transcribed across fixed-$\Sigma_{yy}$ and fixed-$V$ protocols.
In this Letter, we map comprehensively the rheology of dry and wet granular matter, at both fixed $V$ and fixed $\sigma_{yy}$. Our simulation results reveal constitutive curves that are indeed non-monotonic and shear thinning (thickening) at fixed $\sigma_{yy}$ (fixed $V$). We demonstrate a mapping whereby constitutive curves measured at fixed $\sigma_{yy}$ can be used to construct fixed-$V$ counterparts, and vice versa. Performing this on our data, we investigate whether non-monotonicity at fixed $\sigma_{yy}$ implies the same at fixed $V$. We show that this is not so, uncovering instead a signature in the fixed-$\sigma_{yy}$ curves $\phi(I)$ (or $\phi(J)$) that, if present, implies non-monotonic $\sigma_{xy}(\dot{\gamma})$ at fixed $V$. In the context of dry systems, we give the first simulation evidence for shear bands with layer-normals in the gradient direction at fixed $\sigma_{yy}$; and vorticity direction at fixed $V$, consistent with the measured constitutive curves. We furthermore demonstrate a non-monotonic $\mu(J)$ in suspensions, recently suggested experimentally [34]. We show separately that discontinuous shear thickening (DST) can arise at fixed $V$ even in the absence of critical load friction [35, 36], hitherto considered prerequisite.
Simulation.— We simulate the Newtonian dynamics of a packing of spheres of density $\rho$ with bidisperse radii $a$ and $1.4a$ using LAMMPS [37, 38, 39]. Interparticle contacts are modelled as Hookean with stiffness $k$  [^1] We verified that varying $k$ over an order of magnitude does not significantly change any of the results reported., frictional with sliding coefficient $\mu_{p}=0.5$, and damped with restitution coefficient 0.5. For wet systems we also implement short-ranged lubrication forces with viscosity $\eta$, in which the leading term diverges with the inverse of the surface-to-surface distance $h$, truncated when $h<10^{-3}a$ [41].
We impose simple shear of rate ${\dot{\gamma}}=\partial v_{x}/\partial y$ via Lees-Edwards boundaries, with flow, gradient and vorticity directions $x$, $y$ and $z$. Our time unit $\tau_{\rm I}=\sqrt{\rho a^{3}/k}$ and $\tau_{\rm V}=\eta a/k$ in dry and wet systems respectively, giving non-dimensional shear rates $\dot{\gamma}_{d}=\dot{\gamma}\tau_{\rm I}$ and $\dot{\gamma}_{w}=\dot{\gamma}\tau_{\rm V}$. The stress is computed by averaging the tensor product of particle-particle vectors and forces, and rescaled as $\Sigma_{ij}=\sigma_{ij}a/k$, so that $\mu=\Sigma_{xy}/\Sigma_{yy}$, $I=\dot{\gamma}_{d}/\sqrt{\Sigma_{yy}}$ and $J=\dot{\gamma}_{w}/\Sigma_{yy}$. Steady state data are averaged over 30 realisations measured up to 10 strain units beyond the initial transient. Results are independent of our timestep, $\delta t=0.01\tau_{\rm I}$. The periodic box size $L_{x,y,z}\approx 30a$ in Figs. 14, with elongated $L_{y}$ or $L_{z}$ specified in Fig. 23.
To model each of the experimental protocols described above, we perform simulations in two different modes. In one we fix $V$ (and so $\phi$) and ${\dot{\gamma}}$, measuring ${\Sigma_{xy}}$ and ${\Sigma_{yy}}$ in steady state. In the other we fix ${\dot{\gamma}}$ and introduce a set point normal stress $\Sigma_{yy}^{\mathrm{SP}}$, at each timestep updating $V(t+\delta t)=V(t)+\alpha\delta t({\Sigma_{yy}}(t)-\Sigma_{yy}^{\mathrm{SP}})/{\Sigma_{yy}^{\mathrm{SP}}}$  [^2] We set $\alpha=0.1$ having verified that results are insensitive to this choice. and checking that this maintains $\Sigma_{yy}\approx\Sigma_{yy}^{\mathrm{SP}}$ to excellent approximation. We then measure $\phi$ and ${\Sigma_{xy}}$ in steady state.
Dry system.— We start with the fixed-$\Sigma_{yy}$ data. Figures 1(a)-(d) show constitutive curves $\mu(I)$ and $\phi(I)$ obtained for several fixed $\Sigma_{yy}$ by increasing ${\dot{\gamma}_{d}}$ (and so $I=\dot{\gamma}_{d}/\sqrt{\Sigma_{yy}}$) from left to right. Figure 1(e) shows that our control scheme produces a stable ${\Sigma_{yy}}$ over time. Viewed on the coarse scale of Figs. 1(a), (c), the rheology is broadly consistent with that for hard particles, with master curves for $\mu(I)$ and $\phi(I)$ following [22, 25].
Figure 2: Gradient banding in dry system. (a) Strain series of $\dot{\gamma}_{d}$ profile across $y$, measured at fixed global ${\Sigma_{yy}}=0.0016$ and $\bar{\dot{\gamma}}_{d}=10^{-5}$. Binning temporal and spatial $I$ data gives the histogram in (b). From its peaks we obtain $I$ in the low (blue star) and high (red star) shear regions. These are then plotted as $\mu(I)$ in (c). Orange line in (c) represents the homogeneous $\mu(I)$ at $\Sigma_{yy}=0.0016$ from Fig. 1(a). Dashed black lines in (b),(c) indicate imposed global $I$.
However, the separation of $\mu$ and $\phi$ curves on the finer scales of Figs. 1(b),(d), and the anomalous $\phi$ curves in Fig. 1(c) (red lines), show a breakdown of this master scaling for larger $\Sigma_{yy}$, with the flow state not uniquely defined by $I$ but depending separately on $\dot{\gamma}_{d}$ and $\Sigma_{yy}$. Here the particles are slightly compressed due to their finite stiffness $k$. The $\mu(I)$ curves obtained by increasing $\dot{\gamma}_{d}$ at fixed $\Sigma_{yy}$ are then non-monotonic, with a window of $I$ over which $\frac{\partial\mu}{\partial I}|_{\Sigma_{yy}}<0$ (pink boxes, Fig. 1), consistent with an earlier result in 2D [28]. This suggests that flow at an imposed global $I$ in this window will form gradient shear bands of differing $\dot{\gamma}_{d}$ with layer normals along $y$. Meanwhile the $\phi(I)$ data in Fig. 1(d) reveal dilation with increasing $\dot{\gamma}_{d}$ at fixed $\Sigma_{yy}$, with $\frac{\partial\phi}{\partial I}|_{\Sigma_{yy}}<0$.
In this regime, where the flow depends not on $I$ alone but on both $\dot{\gamma}_{d}$ and $\Sigma_{yy}$, one can obtain constitutive curves at fixed $\Sigma_{yy}$ by varying $\dot{\gamma}_{d}$ (as discussed so far) or at fixed $\dot{\gamma}_{d}$ by varying $\Sigma_{yy}$. To explore the latter, we show by the purple line in Fig. 1(d) all data points with a single fixed $\dot{\gamma}_{d}$, now across data for different $\Sigma_{yy}$. Doing so reveals another non-monotonicity, highlighted by the green boxes in Fig. 1, with a window of $I$ for which ${\frac{\partial\phi}{\partial I}}|_{\dot{\gamma}_{d}}>0$, which means that increasing $\Sigma_{yy}$ causes dilation. This will have important implications for the fixed-$V$ rheology, to which we now turn.
Shown in Fig. 1(f) are time-dependent signals $\Sigma_{xy}(t)$ measured at a given $\dot{\gamma}_{d}$ for three $\phi$. Beyond the ${\dot{\gamma}}t\leq 1$ start-up transient, we observe a statistically steady $\Sigma_{xy}$ for $\phi=0.58,~{}0.61$. In contrast, at intermediate $\phi=0.596$ the stress intermittently switches between two apparently metastable values, each sustained for $0.5-2$ strain units. Similar phenomenology was seen in [11], described there as fluctuating rather than bistable behavior 333 Simulations at $\phi=0.596$ and imposed ${\Sigma_{xy}}$ (using an algorithm analogous to that for our fixed-$\Sigma_{yy}$ protocol) invariably led to flow arrest at long times (as reported previously [18, 55]) precluding fixed-$V$ measurement of the unstable region. . To obtain a point (or points) on the stationary constitutive curve from this $\Sigma_{xy}(t)$ signal at each $\dot{\gamma}_{d}$ and $\phi$, we produce a histogram $P(\Sigma_{xy})$ by sampling at intervals of ${\dot{\gamma}}t=0.01$, and fit it to a single Gaussian for low and high $\phi$, or a sum of two Gaussians for bistable cases at intermediate $\phi$. The locations of the maximum (or maxima) of these fits are taken as the time-averaged stress values, shown by solid colored lines in Figs. 1(g),(h).
Figure 3: Vorticity banding in dry system. (a) Strain series of $\Sigma_{xy}$ profile across $z$ measured at fixed global $\phi=0.596$ and $\bar{\Sigma}_{xy}=5\times 10^{-5}$. Binning temporal and spatial $\Sigma_{xy}$ leads to the histogram in (b). From its peaks we obtain the low (pink star) and high (yellow star) $\Sigma_{xy}$ values, plotted in (c) with those obtained under fixed $V$ and under the reconstruction (green and black lines, Fig. 1(g)) at the same $\phi$. Dashed black lines in (b), (c) indicate imposed global $\bar{\Sigma}_{xy}$.
We additionally obtain reconstructed fixed-$V$ constitutive curves from the fixed-$\Sigma_{yy}$ data presented in Figs. 1(a)-(d). Each plotted point therein has known $\mu$, $\phi$ and $I$, with the latter measured at known $\Sigma_{yy}$ and $\dot{\gamma}_{d}$ and with $\Sigma_{xy}=\mu\Sigma_{yy}$. We thus replot these data $\Sigma_{xy}(\dot{\gamma}_{d})$ and $\Sigma_{yy}(\dot{\gamma}_{d})$ as faded colored lines in Fig. 1(g),(h). Here $\Sigma_{yy}(\dot{\gamma}_{d})$ are horizontal lines along which $\phi$ decreases monotonically from left to right. Connecting points on each of these lines that have equal $\phi$, one obtains the solid grey and black contours in Fig. 1(g),(h). For low ${\dot{\gamma}_{d}}$ we observe inertial ${\Sigma_{xy}}=f(\phi){\dot{\gamma}_{d}}^{2}$ rheology [11, 44] for $\phi\leq 0.593$ and quasistatic ${\Sigma_{xy}}=h(\phi)$ for $\phi\geq 0.605$. For large $\dot{\gamma}_{d}$ the data tend towards ‘intermediate’ ${\Sigma_{xy}}\sim{\dot{\gamma}_{d}}^{1/2}$ behavior [11]. This reconstruction produces constitutive curves that are S-shaped and show DST, consistent with the bi-valued constitutive curves obtained for $0.596\leq\phi\leq 0.599$ in the fixed-$V$ simulation. The particle contact number grows towards its isostatic point $z\approx 4$ (for our $\mu_{p}$) with increasing stress at fixed $V$ (data not shown), suggesting a different DST mechanism compared to critical load models in which the isostatic point itself decreases with increasing stress [35, 36].
We now show that these S-shaped $\Sigma_{yy}(\dot{\gamma}_{d})$ curves map directly to a particular feature of $\phi(I)$ in Fig. 1(d). Writing $\phi=f(\dot{\gamma}_{d},\Sigma_{yy})$, one can find the slope of fixed-$V$ curves by setting $d\phi=0$ and using the definition of $I$ to obtain $\frac{d\Sigma_{yy}}{d\dot{\gamma}_{d}}|_{\phi}=(2\Sigma_{yy}/\dot{\gamma}_{d})\left(\frac{\partial f}{\partial I}|_{\Sigma_{yy}}\right)/\left(\frac{\partial f}{\partial I}|_{\dot{\gamma}_{d}}\right)$. Thus if $\frac{\partial\phi}{\partial I}|_{\Sigma_{yy}}<0$ (blue-to-red lines, Fig. 1(d)), then $\frac{\partial\phi}{\partial I}|_{\dot{\gamma}_{d}}>0$ (purple line) implies downward sloping fixed-$V$ curves $\Sigma_{yy}(\dot{\gamma}_{d})$. Indeed, the green boxes in Fig. 1 defined by $\frac{\partial\phi}{\partial I}|_{\dot{\gamma}}>0$ coincide in Fig. 1(g),(h) with $\frac{\partial\Sigma_{yy}}{\partial\dot{\gamma}}|_{\phi}<0$ (and $\frac{\partial\Sigma_{xy}}{\partial\dot{\gamma}}|_{\phi}<0$). The same conclusion may be obtained graphically, by taking a horizontal line through Fig. 1(d) at, say, $\phi=0.598$ (dark grey line). One now finds three intersections with the purple line, giving three stress states for a given $\dot{\gamma}_{d}$ at $\phi=0.598$, implying S-shaped fixed-$V$ constitutive curves. Importantly, the pink and green boxes do not fully overlap, so that non-monotonic $\mu(I)$ (pink) and $\Sigma_{xy}(\dot{\gamma})$ (green) do not imply each other.
So far, we have reported constitutive curves that are non-monotonic and shear thinning at fixed $\Sigma_{yy}$, implying a pre-disposition towards gradient banding; and, in notable contrast, non-monotonic and shear thickening at fixed $\phi$, implying a pre-disposition towards vorticity banding.   [^4] Notably, repeating our fixed-$\Sigma_{yy}$ simulations with $\mu_{p}=0$ leads to monotonic $\mu(I)$ and $\phi(I)$.. We now show that these indeed lead to the formation of bands in simulation boxes large enough to accommodate banding along the relevant axis. (The boxes simulated in Fig. 1 are too small to do so.)
Figure 2 shows a simulation at fixed $\Sigma_{yy}$ in a box elongated along the gradient direction, $L_{y}=1200a$. The stress $\Sigma_{yy}$ is such that the $\mu(I)$ curve for homogeneous shear (in a smaller box) is non-monotonic, Fig. 2(c). A global shear rate $\bar{\dot{\gamma}}_{d}$ is then imposed so that $I$ lies in its negatively sloping part (black star). The time series of the $y$ profile of $\dot{\gamma}_{d}$ then clearly reveals banding, Fig. 2(a). Binning the local $I$ values gives the histogram in Fig. 2(b), with peaks at $I$ values indicated by the blue and red stars in Figs. 2(b),(c). These show that the low shear band is effectively jammed ($I\approx 0$) while the high shear band lies close to the minimum of the homogeneous $\mu(I)$. The stress values remain spatially and temporally uniform (not shown), with $\mu$ in the banded state slightly below that in homogeneous flow, Fig. 2(c). This may be due to non-local effects, which are known to cause deviations from homogeneous rheology that propagate distances $\mathcal{O}(10)a$ from yield planes [46].
Figure 3 reports a fixed-$V$ simulation in a box elongated along the vorticity direction, $L_{z}=1200a$, with $\phi=0.596$ so that $\Sigma_{xy}(\dot{\gamma}_{d})$ is non-monotonic, Fig. 3(c). A set-point global stress $\bar{\Sigma}_{xy}$ (black star) is imposed near its negatively sloping part via an algorithm that dynamically adjusts $\dot{\gamma}_{d}$. The time series of the $z$ profile of $\Sigma_{xy}$, Fig. 3(a), then clearly shows a stress band. This is not however stationary but steadily propagates along $z$, consistent with the balance of normal stresses precluding stationary banding [47]. Binning these $\Sigma_{xy}$ data gives the histogram in Fig. 3(b), with peaks at $\Sigma_{xy}$ values (pink and yellow stars) transcribed to the $\Sigma_{xy}(\dot{\gamma}_{d})$ representation in Fig. 3(c). This finding is contrary to a report in dry systems of a homogeneous jammed state in regions where $\Sigma_{xy}(\dot{\gamma}_{d})$ is expected to be downward sloping [18]. Propagating vorticity bands were however seen (albeit for wet systems with critical load friction) in [48].
Figure 4: Non-monotonic constitutive curves of wet granular material. Shown in (a) and (b) are $\mu(J)$ and $\phi(J)$; in (c) and (d) are the same data along with reconstructed and actual fixed-$V$ data obtained equivalently to those in Fig. 1.
Wet system.— We show finally that wet system constitutive curves have non-monotonicities counterpart to dry systems. Figure 4(a) shows that $\mu(J)$ curves obtained by varying $\dot{\gamma}_{w}$ at fixed $\Sigma_{yy}$ are non-monotonic for a range (albeit limited) of $\Sigma_{yy}$. The $\phi(J)$ curve measured by instead varying $\Sigma_{yy}$ at fixed $\dot{\gamma}_{w}$ is also non-monotonic (purple line, Fig. 4(b)). This then maps to non-monotonic reconstructed S-shaped shear thickening curves $\Sigma_{xy}(\dot{\gamma}_{w})$ and $\Sigma_{yy}(\dot{\gamma}_{w})$ at fixed $V$ in Fig. 4(c),(d), as for dry systems. Simulations performed actually at fixed $V$ in this regime likewise give bi-valued constitutive curves  [^5] The range of $\phi$ for which inaccessible stress states are observed differs between the dry and wet cases, suggesting that the lubrication forces present in the latter system slightly change $\phi_{\rm m}$., with a viscous regime $\Sigma_{xy}\propto\dot{\gamma}_{w}$ for $\phi\leq 0.58$.
For dry systems, we recall that negative slope of the $\mu$ curves at fixed $\Sigma_{yy}$ (pink boxes, Fig. 1) does not automatically imply negative slope of the $\Sigma_{xy}(\dot{\gamma}_{w})$ curves at fixed $V$ (green boxes). This lack of correspondence is even more apparent in wet systems: little overlap is apparent between the pink and green boxes in Fig. 4. Instead, as in dry systems, the presence of negative slope in the constitutive curves $\Sigma_{xy}(\dot{\gamma}_{w})$ at fixed $V$ requires only that the material is dilatant with respect to increases in both $\dot{\gamma}_{w}$ and $\Sigma_{yy}$ under fixed $\Sigma_{yy}$. Grob et al. [18] give a criterion for the existence of such points based on a simple model of additive stress contributions to ${\dot{\gamma}}$.
Conclusion.— We have shown that the constitutive curves of dry and wet granular matter sheared at both fixed $\Sigma_{yy}$ and fixed $V$ are non-monotonic, leading to gradient (vorticity) banding at fixed $\Sigma_{yy}$ (fixed $V$). Mapping between protocols we find that non-monotonicity in $\mu(I)$ (or $\mu(J)$) at fixed $\Sigma_{yy}$ does not imply an S-shaped non-monotonicity in $\Sigma_{xy}(\dot{\gamma}_{d})$ (or $\dot{\gamma}_{w}$) at fixed $V$, or vice-versa. Instead, a non-monotonicity in $\phi(I)$ (or $\phi(J)$) associated with dilatancy under increasing $\dot{\gamma}$ and $\Sigma_{yy}$ maps to non-monotonic $\Sigma_{xy}$ at fixed $V$ and vice versa.
Our simulations also show that DST arises in frictional systems near jamming even in the absence of a critical load model [35, 36], suggesting that finite particle stiffness alone can provide the requisite stress scale. Indeed, contacts of few asperities may render particle softness relevant at contact even for grains of large modulus [50, 51, 52, 53]. Our results pose important questions for constitutive modelling. Models for S-shaped $\Sigma_{xy}(\dot{\gamma})$ exist, but do not predict non-monotonicity in $\mu$ [54, 18, 36]. A phenomenological model encoding flow-induced noise predicts non-monotonic $\mu(I)$ [28], though the demonstration of non-monotonic $\mu(J)$ in this Letter and experimentally [34] calls for an equivalent mechanism to be identified in overdamped systems. Reconciling the rich banding dynamics reported here with a detailed micromechanical description of the flow (accounting also for non-locality [46, 32]) remains an open challenge critical to characterising granular flows in many domains.
Acknowledgements.
The authors thank M. Cates and Y. Jiang for discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 885146) (SMF); and from the Royal Academy of Engineering under the Research Fellowship scheme and from the Leverhulme Trust under Research Project Grant RPG-2022-095 (CN).